# Polynomial invariants for $3$-dimensional Leibniz algebras 111 The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2

### v1

@@
Abstract: For every $3$-dimensional non-Lie Leibniz algebras over complex numbers we describe the algebra of polynomial invariants and the group of automorphisms. As a consequence, we establish that two non-nilpotent $3$-dimensional non-Lie Leibniz algebras can be distinguished by means of the traces of degrees $\leqslant 2$ and the dimensions of groups of automorphisms.
Keywords: Leibniz algebras, polynomial invariants, separating set, generating set.
@@ 1.1 Algebra of invariants
[…]
@@ 1. $\bullet$
    $\mathcal{A}=M_{n}$ is the algebra of all $n\times n$ matrices over $\mathbb{F}$, in case $\mathop{\rm char}{\mathbb{F}}=0$ or $\mathop{\rm char}{\mathbb{F}}>n$. Generators for the algebra of invariants $I_{m}(M_{n})$ were independently described in [36] and [32] in the case of charactristic zero, and in [11] in the case of positive characteristic. Minimal generating sets for $I_{m}(M_{2})$ and $I_{m}(M_{3})$ were obtained in [33, 10, 25, 26, 27] and for $I_{m}(M_{4})$ and $I_{m}(M_{5})$ with small values of $m$, in [39, 12, 9], assuming $\mathop{\rm char}\mathbb{F}=0$.
@@ 2. $\bullet$
    $\mathcal{A}=\mathbf{O}$ is the octonion algebra, in case $\mathop{\rm char}{\mathbb{F}}\neq 2$. Generators for the algebra of invariants $I_{m}(\mathbf{O})$ were constructed in [35] over the field of complex numbers $\mathbb{C}$ and in [40] over an arbitrary infinite field of odd characteristic. In case $\mathop{\rm char}{\mathbb{F}}\neq 2$, a minimal generating set was constructed [28], using the classification of pairs of octonions from [29].
@@ 3. $\bullet$
    $\mathcal{A}={\mathbb{A}}$ is the split Albert algebra, i.e., the exceptional simple Jordan algebra of $3\times 3$ Hermitian matrices over $\mathbf{O}$ with the symmetric multiplication $a\circ b=(ab+ba)/2$, in case $\mathop{\rm char}{\mathbb{F}}=0$ and $m\in\{1,2\}$ (see [16, 31]).
More details about the Artin–Procesi–Iltyakov Equality can be found in [1].
@@ 1.2 Leibniz algebras
Leibniz algebras present a ”non antisymmetric” generalization of Lie algebras. It was at first introduced by A. Bloh [3] and then independently by J.-L. Loday [23]. Recently, they appeared in many geometric and physics applications (see, for example, [4, 22, 21, 37] and references therein). A systematic study of algebraic properties of Leibniz algebras was started by J.-L. Loday and T. Pirashvili [24]. So, several classical theorems from Lie algebras theory have been extended to the Leibniz algebras case; many classification results regarding nilpotent, solvable, etc. Leibniz algebras have been obtained (see, for example, [13, 2, 7, 6, 14, 20, 19, 30, 38] and references therein); for results on algebraic and geometric classifications of low dimensional Leibniz algebras see [17, 18] and references therein.
@@ 1.3 Results
In this section we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. Denote by ${\bf L}_{3}$ the family of all $3$-dimensional non-Lie Leibniz algebras over the field $\mathbb{F}$. The classification of algebras from ${\bf L}_{3}$ modulo the action of the group of automorphisms was independently obtained by I.M. Rikhsiboev, I.S. Rakhimov [34] and by J.M. Casas, M.A. Insua, M. Ladra, S. Ladra [5]. The equivalence of these classifications was pointed out in the Remark at the end of paper [34]. In our paper, we use the classification from [34] (see Theorem 9).
In Section 2, we provide key definitions of generic elements and operator traces. In Section 3, we present an explicit formula for calculating some operator traces (see Proposition 4) and describe the reduction of the general case to multilinear in characteristic zero (see Proposition 6).
Assume $\mathcal{L}\in{\bf L}_{3}$. In Sections 46 we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. In Section 4 we explicitly describe the group $\mathop{\rm Aut}(\mathcal{L})$ of all automorphisms of $\mathcal{L}$ (see Theorem 10). Then in Section 5 we obtain a generating set for the algebra of invariants $I_{m}(\mathcal{L})$ (see Theorem 14). As a consequence, we describe when Artin–Procesi–Iltyakov Equality holds for $\mathcal{L}$ (see Corollary 16). In Corollary 18 we obtain $I_{m}(\mathcal{L})=\mathbb{F}$ for all $m\geqslant 1$ if and only if $\mathcal{L}$ is nilpotent. As a consequence, in Corollary 19 we show that any two non-nilpotent algebras from ${\bf L}_{3}$ can be distinguished by means of
@@ 2. $\bullet$
Note that the problem of distinguishing low-dimensional algebras using their algebraic properties has recently been studied. Specifically, in [8], it was shown that any two-dimensional Jordan algebras over a finite field $\mathbb{F}$ with odd characteristic are isomorphic if and only if they satisfy the same polynomial identities. In [15], it was established that polynomial identities distinguish non-associative two-dimensional Novikov algebras over the complex numbers.
@@ 1.4 Notations
A monomial $w=x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$ has multidegree $\mathop{\rm mdeg}(w)=(\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m}$, where $\delta_{r}$ is the number of letters of $w$ lying in the set $\{x_{r1},\ldots,x_{rn}\}$ and =$\mathbb{N}=\{0,1,2,\ldots\}$. As an example, $\mathop{\rm mdeg}(x_{31}x_{32}x_{21}x_{12})=(1,1,2)$ for $m=3$. For short, we denote the multidegree $(1,\ldots,1)\in\mathbb{N}^{m}$ by $1^{m}$. If $f\in\mathbb{F}[\mathcal{A}^{m}]$ is a linear combination of monomials of the same multidegree ${\underline{d}}$, then we say that $f$ is $\mathbb{N}^{m}$-homogeneous of multidegree ${\underline{d}}$. In other words, we have defined the $\mathbb{N}^{m}$-grading of $\mathbb{F}[\mathcal{A}^{m}]$ by multidegrees. Since $\mathbb{F}$ is infinite, the algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ also has the $\mathbb{N}^{m}$-grading by multidegrees. An $\mathbb{N}$-homogeneous element $f\in\mathbb{F}[\mathcal{A}^{m}]$ of multidegree $(d_{1},\ldots,d_{m})$ with $d_{1},\ldots,d_{m}\in\{0,1\}$ is called multilinear. For short, we denote by $x_{r_{1},i_{1}}\cdots\widehat{x_{r_{l},i_{l}}}\cdots x_{r_{k},i_{k}}$ the monomial $x_{r_{1},i_{1}}\cdots x_{r_{l-1},i_{l-1}}x_{r_{l+1},i_{l+1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$, where $1\leqslant l\leqslant k$. Given a subset $S\subset\mathcal{A}$, we denote by $\mathop{\rm alg}\{S\}$ the subalgebra of $\mathcal{A}$ (without unity in general) generated by $S$. We write $\mathbb{F}^{\times}$ for the set of non-zero elements of $\mathbb{F}$.
@@ 2. $\bullet$
    It is well known that $\mathop{\rm tr}(h)$ is an invariant from $I_{m}(\mathcal{A})$ (for example, see Lemma 3.1 from [1]).
@@ 3.1 Traces
Proposition 4 (Proposition 3.2 of [1]).
@@
Proposition 5 (Proposition 3.3 of [1]).
@@ 3.2 Reduction to multilinear case
${\underline{r}}|l|=j\text{ if and only if }r_{1}+\cdots+r_{j-1}+1\leqslant l\leqslant r_{1}+\cdots+r_{j},$
@@
Lemma 7.
@@
The following lemma follows immediately from a partial case of Theorem 7.6 from [1] for the algebra ${\mathbf{A}}_{2}$.
Lemma 8.
@@
[…]
@@ 4 Automorphisms of $3$-dimensional Leibniz algebras
$\mathcal{L}_{1}$ :::: $\left(\begin{array}[]{ccc}0&0&-2e_{1}\\ 0&e_{1}&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{2}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&\lambda e_{1}\\ 0&0&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{3}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-e_{2}\\ 0&e_{2}&e_{1}\\ \end{array}\right),$
$\mathcal{L}_{4}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&e_{1}\\ 0&0&\lambda e_{1}\\ \end{array}\right),$ $\mathcal{L}_{5}$ :::: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{6}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&e_{1}\\ 0&0&0\\ \end{array}\right),$
$\mathcal{L}_{7}^{\lambda}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&\lambda e_{1}+e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{8}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{9}$ :::: $\left(\begin{array}[]{ccc}0&0&e_{1}+e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$

### v2

# Polynomial invariants for $3$-dimensional Leibniz algebras  [^1] The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2

@@
Abstract: For each $3$-dimensional non-Lie Leibniz algebra over the complex numbers, we describe the algebra of polynomial invariants and determine its group of automorphisms. As a consequence, we establish that any two non-nilpotent $3$-dimensional non-Lie Leibniz algebras can be distinguished by the traces of degrees $\leqslant 2$ and by the dimensions of their automorphism groups.
Keywords: Leibniz algebras, polynomial invariants, generating set.
@@ 1.1 Algebra of invariants
$\mathbb{F}[\mathcal{A}^{m}]=\mathbb{F}[x_{ri}\,|\,1\leqslant r\leqslant m,\;1\leqslant i\leqslant n].$
@@ 1. $\bullet$
    $\mathcal{A}=M_{n}$ is the algebra of all $n\times n$ matrices over $\mathbb{F}$, in case $\mathop{\rm char}{\mathbb{F}}=0$ or $\mathop{\rm char}{\mathbb{F}}>n$. Generators for the algebra of invariants $I_{m}(M_{n})$ were independently described in [Sibirskii68] and [Procesi76] in the case of characteristic zero, and in [Donkin92a] in the case of positive characteristic. Minimal generating sets for $I_{m}(M_{2})$ and $I_{m}(M_{3})$ were obtained in [Procesi84, DKZ02, Lopatin_Comm1, Lopatin_Sib, Lopatin_Comm2] and for $I_{m}(M_{4})$ and $I_{m}(M_{5})$ with small values of $m$, in [Teranishi86, Drensky_Sadikova_4x4, Djokovic07], assuming $\mathop{\rm char}\mathbb{F}=0$.
@@ 2. $\bullet$
    $\mathcal{A}=\mathbf{O}$ is the octonion algebra, in case $\mathop{\rm char}{\mathbb{F}}\neq 2$. Generators for the algebra of invariants $I_{m}(\mathbf{O})$ were constructed in [schwarz1988] over the field of complex numbers $\mathbb{C}$ and in [zubkov2018] over an arbitrary infinite field of odd characteristic. In case $\mathop{\rm char}{\mathbb{F}}\neq 2$, a minimal generating set was constructed [LZ_2], using the classification of pairs of octonions from [LZ_1].
@@ 3. $\bullet$
    $\mathcal{A}={\mathbb{A}}$ is the split Albert algebra, i.e., the exceptional simple Jordan algebra of $3\times 3$ Hermitian matrices over $\mathbf{O}$ with the symmetric multiplication $a\circ b=(ab+ba)/2$, in case $\mathop{\rm char}{\mathbb{F}}=0$ and $m\in\{1,2\}$ (see [Iltyakov_1995, Polikarpov_1991]).
  4. $\bullet$
    $\mathcal{A}$ is a two-dimensional simple algebra with a non-trivial automorphism group, in case $\mathop{\rm char}{\mathbb{F}}=0$ (see [Alvarez_Lopatin_2025]).
More details about the Artin–Procesi–Iltyakov Equality can be found in [Alvarez_Lopatin_2025].
@@ 1.2 Leibniz algebras
Leibniz algebras present a ”non antisymmetric” generalization of Lie algebras. It was at first introduced by A. Bloh [Bloh_1965] and then independently by J.-L. Loday [L93a]. Recently, they appeared in many geometric and physics applications (see, for example, [bonez, kotov20, KW01, strow20] and references therein). A systematic study of algebraic properties of Leibniz algebras was started by J.-L. Loday and T. Pirashvili [L93]. So, several classical theorems from Lie algebras theory have been extended to the Leibniz algebras case; many classification results regarding nilpotent, solvable, etc. Leibniz algebras have been obtained (see, for example, [FKS_2025, AKS, deMello_Souza_2023, CK, FW, KM, kky22, MS, TX] and references therein); for results on algebraic and geometric classifications of low dimensional Leibniz algebras see [k23, KKP] and references therein.
@@ 1.3 Results
In this section, we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. Denote by ${\bf L}_{3}$ the family of all $3$-dimensional non-Lie Leibniz algebras over the field $\mathbb{F}$. The classification of algebras from ${\bf L}_{3}$ modulo the action of the automorphism group was independently obtained by:
  1. $\bullet$
    J.M. Casas, M.A. Insua, M. Ladra, and S. Ladra [Casas_Insua_Ladra_Ladra_2012_LAA], by means of a computer program;
  2. $\bullet$
    I.M. Rikhsiboev and I.S. Rakhimov [Rikhsiboev_Rakhimov_2012_AIP], without the use of a computer. Unfortunately, this classification is not complete (see the proof of Theorem 9 for details).
In our paper, we complete the classification given in [Rikhsiboev_Rakhimov_2012_AIP] in Theorem 9, and we prove its equivalence with the classification obtained in [Casas_Insua_Ladra_Ladra_2012_LAA] (see again the proof of Theorem 9).
In Section 2, we provide the key definitions of generic elements and operator traces. In Section 3, we present explicit formulas for calculating certain operator traces (see Proposition 4) and describe the reduction of the general case to the multilinear case in characteristic zero (see Proposition 6).
Assume $\mathcal{L}\in{\bf L}_{3}$. In Sections 46 we assume that $\mathbb{F}=\mathbb{C}$ is the field of complex numbers. In Section 4 we explicitly describe the group $\mathop{\rm Aut}(\mathcal{L})$ of all automorphisms of $\mathcal{L}$ (see Theorem 10). Then in Section 5 we obtain a generating set for the algebra of invariants $I_{m}(\mathcal{L})$ (see Theorem 14). As a consequence, we describe when the Artin–Procesi–Iltyakov Equality holds for $\mathcal{L}$ (see Corollary 16). In Corollary 18 we obtain $I_{m}(\mathcal{L})=\mathbb{F}$ for all $m\geqslant 1$ if and only if $\mathcal{L}$ is nilpotent. As a consequence, in Corollary 19 we show that any two non-nilpotent algebras from ${\bf L}_{3}$ can be distinguished by means of
@@ 2. $\bullet$
Note that the problem of distinguishing low-dimensional algebras using their algebraic properties has recently been studied. Specifically, in [diniz2024isomorphism], it was shown that any two-dimensional Jordan algebras over a finite field $\mathbb{F}$ with odd characteristic are isomorphic if and only if they satisfy the same polynomial identities. In [PIs_Novikov_dim2_2025], it was established that polynomial identities distinguish non-associative two-dimensional Novikov algebras over the complex numbers.
@@ 1.4 Notations
A monomial $w=x_{r_{1},i_{1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$ has multidegree $\mathop{\rm mdeg}(w)=(\delta_{1},\ldots,\delta_{m})\in\mathbb{N}^{m}$, where $\delta_{r}$ is the number of letters of $w$ lying in the set $\{x_{r1},\ldots,x_{rn}\}$ and $\mathbb{N}=\{0,1,2,\ldots\}$. As an example, $\mathop{\rm mdeg}(x_{31}x_{32}x_{21}x_{12})=(1,1,2)$ for $m=3$. For short, we denote the multidegree $(1,\ldots,1)\in\mathbb{N}^{m}$ by $1^{m}$. If $f\in\mathbb{F}[\mathcal{A}^{m}]$ is a linear combination of monomials of the same multidegree $\underline{d}$}, then we say that $f$ is $\mathbb{N}^{m}$-homogeneous of multidegree $\underline{d}$}. In other words, we have defined the $\mathbb{N}^{m}$-grading of $\mathbb{F}[\mathcal{A}^{m}]$ by multidegrees. Since $\mathbb{F}$ is infinite, the algebra of invariants $\mathbb{F}[\mathcal{A}^{m}]^{G}$ also has the $\mathbb{N}^{m}$-grading by multidegrees. An $\mathbb{N}$-homogeneous element $f\in\mathbb{F}[\mathcal{A}^{m}]$ of multidegree $(d_{1},\ldots,d_{m})$ with $d_{1},\ldots,d_{m}\in\{0,1\}$ is called multilinear. For short, we denote by $x_{r_{1},i_{1}}\cdots\widehat{x_{r_{l},i_{l}}}\cdots x_{r_{k},i_{k}}$ the monomial $x_{r_{1},i_{1}}\cdots x_{r_{l-1},i_{l-1}}x_{r_{l+1},i_{l+1}}\cdots x_{r_{k},i_{k}}$ from $\mathbb{F}[\mathcal{A}^{m}]$, where $1\leqslant l\leqslant k$. Given a subset $S\subset\mathcal{A}$, we denote by $\mathop{\rm alg}\{S\}$ the subalgebra of $\mathcal{A}$ (without unity in general) generated by $S$. We write $\mathbb{F}^{\times}$ for the set of non-zero elements of $\mathbb{F}$.
@@ 2. $\bullet$
    It is well known that $\mathop{\rm tr}(h)$ is an invariant from $I_{m}(\mathcal{A})$ (for example, see Lemma 3.1 from [Alvarez_Lopatin_2025]).
@@ 3.1 Traces
Proposition 4 (Proposition 3.2 of [Alvarez_Lopatin_2025]).
@@
Proposition 5 (Proposition 3.3 of [Alvarez_Lopatin_2025]).
@@ 3.2 Reduction to multilinear case
${\underline{r}}|l|=j\text{ if and only if }\sum_{1\leqslant i\leqslant j-1}r_{i}+1\leqslant l\leqslant\sum_{1\leqslant i\leqslant j}r_{i},$
@@
Lemma 7 (cf. Theorem 5.14 of [Bruns_Gubeladze_book_2009]).
@@
The following lemma follows immediately from a partial case of Theorem 7.6 from [Alvarez_Lopatin_2025] for the algebra ${\mathbf{A}}_{2}$.
Lemma 8 (see Theorem 7.6 of [Alvarez_Lopatin_2025]).
@@
where $\alpha\in\mathbb{F}$. As in Section 1.1, the coordinate ring $\mathbb{F}[\mathcal{V}^{m}]=\mathbb{F}[x_{r1},x_{r2}\,|\,1\leqslant r\leqslant m]$ of $\mathcal{V}^{m}$ is a $G$-module. Then the algebra of invariants $\mathbb{F}[\mathcal{V}^{m}]^{G}$ is generated by
@@ 4 Automorphisms of $3$-dimensional Leibniz algebras
$\mathcal{L}_{1}$ :: $\left(\begin{array}[]{ccc}0&0&-2e_{1}\\ 0&e_{1}&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $\left(\begin{array}[]{ccc}0&0&\lambda e_{1}\\ 0&0&-e_{2}\\ 0&e_{2}&0\\ \end{array}\right),$ $\mathcal{L}_{3}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&-e_{2}\\ 0&e_{2}&e_{1}\\ \end{array}\right),$
$\mathcal{L}_{4}^{\lambda}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&e_{1}\\ 0&0&\lambda e_{1}\\ \end{array}\right),$ $\mathcal{L}_{5}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&e_{1}&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{6}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&e_{1}\\ 0&0&0\\ \end{array}\right),$
$\mathcal{L}_{7}^{\lambda}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&\lambda e_{1}+e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{8}$ :: $\left(\begin{array}[]{ccc}0&0&e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$ $\mathcal{L}_{9}$ :: $\left(\begin{array}[]{ccc}0&0&e_{1}+e_{2}\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right),$
$\mathcal{L}_{10}$ :: $\left(\begin{array}[]{ccc}0&0&e_{1}\\ 0&0&e_{2}\\ 0&0&0\\ \end{array}\right),$ $\mathcal{L}_{11}$ :: $\left(\begin{array}[]{ccc}0&0&0\\ 0&0&0\\ 0&0&e_{1}\\ \end{array}\right).$
@@
Proof.
A classification of 3-dimensional non-Lie Leibniz algebras over complex numbers was obtained by Rikhsiboev, Rakhimov in Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP], where it was claimed that any algebra from ${\bf L}_{3}$ belongs to the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{6}$, $\mathcal{L}_{7}^{\lambda}$, $\mathcal{L}_{8}$, $\mathcal{L}_{9}$ with $\lambda\in\mathbb{F}$. Unfortunately, this classification is not complete:
  1. $\bullet$
    The algebra $\mathcal{L}_{2}^{0}$ is a Lie algebra, therefore, we should assume that $\lambda$ is non-zero in case $\mathcal{L}_{2}^{\lambda}$.
  2. $\bullet$
    On page 361 of the proof of Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP] it was claimed that $\mathcal{L}_{10}$ is isomorphic to $\mathcal{L}_{6}$, which is not the case. Actually, as it follows from Corollary 19 (see below), $\mathcal{L}_{10}$ is not isomorphic to any other algebra from the formulation of Theorem 9.
  3. $\bullet$
    The algebra $\mathcal{L}_{11}$ was missed in Case 2.1 of the proof of Theorem 2 of [Rikhsiboev_Rakhimov_2012_AIP]. To show that $\mathcal{L}_{11}$ is not isomorphic to any other algebra from the formulation of Theorem 9, we note that the only commutative algebras from the formulation of Theorem 9 are $\mathcal{L}_{5}$ and $\mathcal{L}_{11}$, but $\dim{\rm Ann}^{\rm R}(\mathcal{L}_{5})=1$ and $\dim{\rm Ann}^{\rm R}(\mathcal{L}_{11})=2$, where the right annihilator of an algebra $\mathcal{A}$ is Ann^R(A) ={a∈A —  ba=0 for all b∈A}.
To complete the proof, we will prove that the classification from Theorem 9 is equivalent to the computer-based classification of 3-dimensional non-Lie Leibniz algebras over complex numbers by J.M. Casas, M.A. Insua, M. Ladra, S. Ladra [Casas_Insua_Ladra_Ladra_2012_LAA]. For $\delta\in\mathbb{F}$, we denote by
$\mathcal{L}_{2a}^{\delta},\;\mathcal{L}_{2b},\;\mathcal{L}_{2c},\;\mathcal{L}_{2d},\;\mathcal{L}_{2e}^{\delta}\;(\delta\neq 0),\;\mathcal{L}_{2f},\;\mathcal{L}_{2g},\;\mathcal{L}_{3a}^{\delta}\;(\delta\in\widetilde{\mathbb{F}}),\;\mathcal{L}_{3b},\;\mathcal{L}_{3c},\;\mathcal{L}_{3d}$ (4.1)
the 3-dimensional Leibniz algebras from page 3752 of [Casas_Insua_Ladra_Ladra_2012_LAA], where
$\widetilde{\mathbb{F}}=\{\alpha\in\mathbb{C}\,|\,|\alpha|>1\}\cup\{\alpha\in\mathbb{C}\,|\,|\alpha|=1,\;0\leqslant{\rm arg}(\alpha)\leqslant\pi\}.$
[…]
  1. $\bullet$
    for each line, the algebra from the first column is isomorphic to the algebra from the second column;
  2. $\bullet$
    the tableau of multiplication of the algebra from the first column considered with respect to the ordered basis from the third column is equal to the tableau of multiplication of the algebra from the second column.
$\mathcal{L}_{1}$ $\mathcal{L}_{2g}$ $\{e_{1},e_{1}+e_{2},-e_{2}-e_{3}\}$
$\mathcal{L}_{2}^{\lambda\neq 0}$ $\mathcal{L}_{2e}^{-\lambda}$ $\{e_{1},e_{2},-e_{3}\}$
$\mathcal{L}_{3}$ $\mathcal{L}_{2f}$ $\{e_{1},e_{2},-e_{3}\}$
$\mathcal{L}_{4}^{\lambda\neq 0}$ $\mathcal{L}_{2a}^{\lambda}$ $\{e_{1},e_{3},e_{2}\}$
$\mathcal{L}_{4}^{0}$ $\mathcal{L}_{2a}^{0}$ $\{e_{1},e_{3},e_{2}\}$
$\mathcal{L}_{5}$ $\mathcal{L}_{2c}$ $\{e_{1},e_{2},e_{3}\}$
$\mathcal{L}_{6}$ $\mathcal{L}_{3a}^{-1}$ $\{e_{1}+e_{2},e_{2}-e_{1},-e_{3}\}$
$\mathcal{L}_{7}^{\lambda\neq-\frac{1}{4},0}$ $\mathcal{L}_{3a}^{\lambda^{\prime}}$ $\{\xi_{1}e_{1}+e_{2},\xi_{2}e_{1}+e_{2},\nu e_{3}\}$
$\mathcal{L}_{7}^{-\frac{1}{4}}$ $\mathcal{L}_{3b}$ $\{e_{2},-\frac{e_{1}}{2}+e_{2},2e_{3}\}$
$\mathcal{L}_{7}^{0}$ $\mathcal{L}_{2d}$ $\{e_{2},e_{1}-e_{2},e_{3}\}$
$\mathcal{L}_{8}$ $\mathcal{L}_{3c}$ $\{e_{1},e_{2},e_{3}\}$
$\mathcal{L}_{9}$ $\mathcal{L}_{3d}$ $\{2e_{1}+e_{2},2e_{1}+2e_{2},e_{1}+e_{3}\}$
$\mathcal{L}_{10}$ $\mathcal{L}_{3a}^{1}$ $\{e_{1},e_{2},e_{3}\}$
$\mathcal{L}_{11}$ $\mathcal{L}_{2b}$ $\{e_{1},e_{2},e_{3}\}$
Here,
  1. $\bullet$
[…]
  2. $\bullet$
    $\nu=\frac{2\xi_{1}+1}{\xi_{1}-2\lambda}$.
[…]
$\xi_{1}\xi_{2}=-\lambda\;\text{ and }\;\xi_{1}+\xi_{2}=-1.$ (4.2)
Since $\lambda\neq-1/4$, we have
$\xi_{1}\neq 2\lambda\;\text{ and }\;\xi_{2}\neq 2\lambda.$ (4.3)
[…]
Formulas (4.3) imply that $\nu$ is well-defined and $\nu\neq 0$. Hence,
$v_{1}=\xi_{1}e_{1}+e_{2},\;\;v_{2}=\xi_{2}e_{1}+e_{2},\;\;v_{3}=\nu e_{3}$
is a basis for $\mathcal{L}$.
Compute the tableau of multiplication for $\mathcal{L}$ with respect to the ordered basis $\{v_{1},v_{2},v_{3}\}$:
  1. $\bullet$
    Using formulas (4.2) it is not difficult to see that $\nu\lambda=\lambda^{\prime}\xi_{1}$ and $\nu(\xi_{1}+1)=\lambda^{\prime}$. Hence, v_1v_3=ξ_1νe_2 + ν(λe_1+e_2) = νλe_1 + ν(ξ_1+1) e_2 = λξ_1 e_1 + λ’ e_2 = λ’ v_1.
  2. $\bullet$
    Using formulas (4.2) it is not difficult to see that $\nu\lambda=\xi_{2}$ and $\nu(\xi_{2}+1)=1$. Hence, v_2v_3=ξ_2νe_2 + ν(λe_1+e_2) = νλe_1 + ν(ξ_2+1) e_2 = ξ_2 e_1 + e_2 = v_2.
  3. $\bullet$
    others products $v_{i}v_{j}$ are zero.
Therefore, the tableau of multiplication for $\mathcal{L}$ with respect to the ordered basis $\{v_{1},v_{2},v_{3}\}$ is
$\left(\begin{array}[]{ccc}0&0&\lambda^{\prime}v_{1}\\ 0&0&v_{2}\\ 0&0&0\\ \end{array}\right).$
Hence, $\mathcal{L}\simeq\mathcal{L}_{3a}^{\lambda^{\prime}}$. The following claim, which can be proven by straightforward computations, completes the proof:
[…]
where $\xi_{1},\xi_{2}\in\mathbb{F}$ are pairwise different roots of the equation $x^{2}+x-\lambda=0$.